# FUSION: Forecast-Embedded Agent Scheduling with Service Incentive Optimization over Distributed Air-Ground Edge Networks

### v1

@@ IV-C3 Offline Auction-based Incentive-Compatible Contracting (Off-AIC2)
Step 2. ES–AgP matching and route planning via eACO-VRP. Given the pivotal ranks $(I^{\ast},K^{\ast})$, only the top-$I^{\ast}$ ESs and top-$K^{\ast}$ AgPs proceed to the matching stage. We initialize the set of unmatched ESs as $\mathcal{U}^{\mathrm{free}}=\{1,\dots,I^{\ast}\}$ and iterate over AgP sellers $s_{k}$ in the order of $\mathcal{L}^{\mathrm{AgP}}$. For each $s_{k}$ with remaining capacity $C_{k}^{\mathrm{V}}>0$ and $\mathcal{U}^{\mathrm{free}}\neq\emptyset$, we construct a candidate ES set $\mathcal{C}_{k}=\big\{s_{i}\in\mathcal{U}^{\mathrm{free}}\,\big|\,U_{i,k}>0,\;\mathrm{bid}_{i,k}\geq\overline{\mathrm{bid}}_{i}\big\}$, where $U_{i,k}=\mathrm{bid}_{i,k}-\mathrm{ask}_{k}$ denotes the marginal utility of serving ES $s_{i}$ with AgP $v_{k}$. Then, $v_{k}$ invokes the eACO-VRP (see Alg. 1), which explicitly incorporates the individual information of ESs and AgPs, such as ES demand volumes and time windows, travel time and energy cost between ESs, and the payment that $v_{k}$ can receive, to jointly determine: (i) which ESs in $\mathcal{C}_{k}$ to serve within its residual capacity, and (ii) in what visiting order they should be placed along a feasible route. By maximizing the AgP’s utility along the constructed route, Alg. 1 outputs a selected ES set $\mathcal{K}_{k}^{\ast}\subseteq\mathcal{C}_{k}$ and the associated pre-planned path for AgP $v_{k}$. For every $s_{i}\in\mathcal{K}_{k}^{\ast}$, we set $x_{i,k}^{\mathrm{Off}}=1$ and remove $s_{i}$ from $\mathcal{U}^{\mathrm{free}}$, ensuring that each ES is matched to at most one AgP. Stacking the decisions across all AgPs yields the assignment matrix $\bm{X}^{\ast}$. Moreover, by enforcing $\mathrm{bid}_{i,k}\geq\overline{\mathrm{bid}}_{i}$ for all successful pairs $(s_{i},s_{k})$, the matching stage respects ES-side individual rationality.
Step 3. Pricing and contract finalization via binary search. Once $\bm{X}^{\ast}$ is obtained, Off-AIC2 determines ES payments and AgP rewards via a monotone binary-search procedure designed to approximate truthfulness and budget balance. For each matched ES buyer $s_{i}$ (with $\sum_{k}x^{\mathrm{Off}}_{i,k}=1$), we treat $\overline{\mathrm{bid}}_{i}$ as an effective bid parameter and iteratively search within $\overline{\mathrm{bid}}_{I^{\ast}+1},\,\overline{\mathrm{bid}}_{i}]$ to find the minimal critical payment $p_{i}^{\mathrm{b}\ast}$ that keeps $s_{i}$ in the winning set when Phase 2 is re-executed under this effective bid. Symmetrically, for each matched AgP seller $s_{k}$ (with $\sum_{i}x^{\mathrm{Off}}_{i,k}\geq 1$), we adjust its effective ask within $\mathrm{ask}_{K^{\ast}+1},\,\mathrm{ask}_{k}]$ and re-run Phase 2 to determine the maximal acceptable reward $r_{k}^{\mathrm{s}\ast}$ that still preserves its winning status. The resulting settlement vectors $(\bm{p}^{\mathrm{b}\ast},\bm{r}^{\mathrm{s}\ast})$ finalize the contracts between ESs and AgPs, leading to fair and incentive-aware trading outcomes under the proposed Off-AIC2.
IV-D Explainability and Effectiveness Analysis (e.g., From the View of Auction Properties)
Definition 1.
(Individual Rationality) An auction is said to be individually rational if the payment made by any winning ES (to its matched AgP) does not exceed its bid, and the revenue obtained by any winning AgP is no less than its asked price.
Definition 2.
(Truthfulness) An auction is truthful (or incentive-compatible) if all ESs and AgPs report their bids and asked prices according to their true private valuations; that is, no participant can benefit from misreporting its information.
Definition 3.
(Budget Balance) An auction satisfies budget balance if the total payment collected from all winning ESs is no less than the total payment made to all winning AgPs.
The proposed offline stage is shown to meet all the above properties, with detailed theoretical analysis and proofs moved to Appx. C due to space limitations.
@@
V Online Stage: Potential Game for Service Scheduling over Human-Machine Ecosystem
@@
Facing real-time demand from both HuUs and MaUs covered by ESs, we now turn to an online stage, for which we consider two types of players: (i) SPs (e.g., ESs and AgPs), collected as $\mathcal{S}^{\mathrm{On}}=\{s^{\mathrm{On}}_{1},\ldots,s^{\mathrm{On}}_{\mathbbm{i}},\ldots,s^{\mathrm{On}}_{|\mathcal{S}^{\mathrm{On}}|}\}$; and (ii) SDs (e.g., HuUs and MaUs), collected as $\mathcal{U}^{\mathrm{On}}=\{u^{\mathrm{On}}_{1},\ldots,u^{\mathrm{On}}_{\mathbbm{j}},\ldots,u^{\mathrm{On}}_{|\mathcal{U}^{\mathrm{On}}|}\}$. In this stage, each SD strategically decides whether to execute tasks locally or offload them to one of the available SPs, while SPs allocate their limited computing capacities among the associated tasks. Such mutually coupled decisions are modeled as a game-theoretic scheduling problem.
To represent the decision of $u^{\mathrm{On}}_{\mathbbm{j}}$, we define the binary indicator vector
$\bm{X}^{\mathrm{On}}_{\mathbbm{j}}=\big(x^{\mathrm{On}}_{\mathbbm{j}},\,x^{\mathrm{On}}_{\mathbbm{j},1},\,\ldots,\,x^{\mathrm{On}}_{\mathbbm{j},\mathbbm{i}},\,\ldots,\,x^{\mathrm{On}}_{\mathbbm{j},|\mathcal{S}^{\mathrm{On}}|}\big)\in\{0,1\}^{\,1+|\mathcal{S}^{\mathrm{On}}|},$ (9)
where $x^{\mathrm{On}}_{\mathbbm{j}}=1$ indicates local execution at $u^{\mathrm{On}}_{\mathbbm{j}}$, and $x^{\mathrm{On}}_{\mathbbm{j},\mathbbm{i}}=1$ indicates offloading to provider $s^{\mathrm{On}}_{\mathbbm{i}}$, with $x^{\mathrm{On}}_{\mathbbm{j}}\in\{0,1\}$ and $x^{\mathrm{On}}_{\mathbbm{j},\mathbbm{i}}\in\{0,1\}$ for all $u^{\mathrm{On}}_{\mathbbm{j}}$.
Since each task is either executed locally or offloaded to exactly one SP, we have
$x^{\mathrm{On}}_{\mathbbm{j}}+\sum_{\mathbbm{i}=1}^{|\mathcal{S}^{\mathrm{On}}|}x^{\mathrm{On}}_{\mathbbm{j},\mathbbm{i}}=1,\quad\forall\,\mathbbm{j}\in\{1,\ldots,|\mathcal{U}^{\mathrm{On}}|\}.$ (10)
@@ V-A1 Delay and Congestion Model
We consider SP set $\mathcal{S}^{\mathrm{On}}$ and SD set $\mathcal{U}^{\mathrm{On}}=\bm{\mathcal{U}}^{\mathrm{H}}\cup\bm{\mathcal{U}}^{\mathrm{A}}$, where $\bm{\mathcal{U}}^{\mathrm{H}}$ and $\bm{\mathcal{U}}^{\mathrm{A}}$ denote the sets of HuUs and MaUs, respectively. Each SP $s^{\mathrm{On}}_{\mathbbm{i}}$ has a finite computing capacity $f^{\mathrm{On}}_{\mathbbm{i}}>0$ (CPU cycles/s). When a subset of SDs $\mathcal{U}_{\mathbbm{i}}=\{u^{\mathrm{On}}_{\mathbbm{v}}\in\mathcal{U}^{\mathrm{On}}:x^{\mathrm{On}}_{\mathbbm{v},\mathbbm{i}}=1\}$ simultaneously offload their workload to $s^{\mathrm{On}}_{\mathbbm{i}}$, we adopt a weighted processor sharing (WPS) rule[13] to capture computing congestion and differentiated service levels.
Each SD $u^{\mathrm{On}}_{\mathbbm{j}}$ associated with $s^{\mathrm{On}}_{\mathbbm{i}}$ is assigned a positive weight $\gamma_{\mathbbm{j},\mathbbm{i}}>0$ that reflects service priority, contract tier, or fairness quota. Under WPS, the effective computing rate that $u^{\mathrm{On}}{\mathbbm{j}}$ receives from $s^{\mathrm{On}}{\mathbbm{i}}$ is $\hat{f}_{\mathbbm{j},\mathbbm{i}}=f^{\mathrm{On}}_{\mathbbm{i}}\frac{\gamma_{\mathbbm{j},\mathbbm{i}}}{\sum_{u^{\mathrm{On}}_{\mathbbm{v}}\in\mathcal{U}^{\mathrm{On}}_{\mathbbm{i}}}\gamma_{\mathbbm{v},\mathbbm{i}}}$, and the corresponding edge-side computing delay is
@@
where $r_{\mathbbm{j}}$ denotes the computing workload (e.g., CPU cycles needed) of $u^{\mathrm{On}}_{\mathbbm{j}}$. This WPS model compactly captures how the delay of each individual SD grows with the aggregate weighted load at the SP.
To facilitate analysis, we then define the effective congestion load at $s^{\mathrm{On}}_{\mathbbm{i}}$ as $y_{\mathbbm{i}}=\sum_{u^{\mathrm{On}}_{\mathbbm{v}}\in\mathcal{U}^{\mathrm{On}}_{\mathbbm{i}}}\gamma_{\mathbbm{v},\mathbbm{i}}$, so that (11) can be written in a factorized form as
$t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}=\theta_{\mathbbm{j},\mathbbm{i}}\;\tau_{\mathbbm{i}}(y_{\mathbbm{i}}),\quad\theta_{\mathbbm{j},\mathbbm{i}}=\frac{r_{\mathbbm{j}}}{\gamma_{\mathbbm{j},\mathbbm{i}}},\quad\tau_{\mathbbm{i}}(y)=\frac{y}{f^{\mathrm{On}}_{\mathbbm{i}}}.$ (12)
where $\theta_{\mathbbm{j},\mathbbm{i}}$ is a user-specific load factor, whereas $\tau_{\mathbbm{i}}(y)$ captures the server-side congestion effect as a function of the aggregate load. This separable structure will be instrumental for constructing an ordinal potential in the subsequent analysis. In addition, when $u^{\mathrm{On}}_{\mathbbm{j}}$ offloads to $s^{\mathrm{On}}_{\mathbbm{i}}$, the uplink transmission delay is given by $t^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}=\frac{d_{\mathbbm{j}}}{W\log_{2}\!\big(1+e^{\mathrm{tx}}_{\mathbbm{j}}\,\Gamma_{\mathbbm{j},\mathbbm{i}}\big)}$, where $d_{\mathbbm{j}}$ is the input data size, $W$ is the communication bandwidth, $e^{\mathrm{tx}}_{\mathbbm{j}}$ is the transmit power, and $\Gamma_{\mathbbm{j},\mathbbm{i}}$ is the SNR-related channel gain between $u^{\mathrm{On}}_{\mathbbm{j}}$ and $s^{\mathrm{On}}_{\mathbbm{i}}$. The total edge-side latency experienced by $u^{\mathrm{On}}_{\mathbbm{j}}$ when uploading its workload to $s^{\mathrm{On}}_{\mathbbm{i}}$ is then
@@
V-A2 Utility of HuUs
Consider a HuU $u^{\mathrm{On}}_{\mathbbm{j}}\in\bm{\mathcal{U}}^{\mathrm{H}}$ with local CPU frequency $f^{\mathrm{loc}}_{\mathbbm{j}}$ and local energy coefficient $e^{\mathrm{loc}}_{\mathbbm{j}}$. The local execution time and energy consumption of its task are $t^{\mathrm{loc}}_{\mathbbm{j}}=\frac{r_{\mathbbm{j}}}{f^{\mathrm{loc}}_{\mathbbm{j}}}$ and $E^{\mathrm{loc}}_{\mathbbm{j}}=\frac{r_{\mathbbm{j}}\,e^{\mathrm{loc}}_{\mathbbm{j}}}{f^{\mathrm{loc}}_{\mathbbm{j}}}$. By transferring its task to $s^{\mathrm{On}}_{\mathbbm{i}}$, the HuU achieves a latency saving $t^{\mathrm{save}}_{\mathbbm{j},\mathbbm{i}}=t^{\mathrm{loc}}_{\mathbbm{j}}-\big(t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}+t^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}\big)$, and an energy saving $c^{\mathrm{save}}_{\mathbbm{j},\mathbbm{i}}=E^{\mathrm{loc}}_{\mathbbm{j}}-E^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}$, where $E^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}$ denotes the uplink transmission energy.
We introduce non-negative valuation coefficients $\mathbb{V}^{\mathrm{H}}_{t}$ and $\mathbb{V}^{\mathrm{H}}_{e}$ to capture HuU sensitivities to latency and energy, respectively. The HuU’s perceived valuation of being served by $s^{\mathrm{On}}_{\mathbbm{i}}$ is modeled as $v^{\mathrm{H}}_{\mathbbm{j},\mathbbm{i}}=\mathbb{V}^{\mathrm{H}}_{t}\,t^{\mathrm{save}}_{\mathbbm{j},\mathbbm{i}}+\mathbb{V}^{\mathrm{H}}_{e}\,c^{\mathrm{save}}_{\mathbbm{j},\mathbbm{i}}$. If $s^{\mathrm{On}}_{\mathbbm{i}}$ charges a price $p_{\mathbbm{j},\mathbbm{i}}$ for serving $u^{\mathrm{On}}_{\mathbbm{j}}$, the HuU’s utility becomes

### v2

@@ IV-C3 Offline Auction-based Incentive-Compatible Contracting (Off-AIC2)
For such an AP $v_{k}$, we invoke the eACO-VRP routine (Sec. IV-C2), which considers ES demand volumes over time, travel time and energy cost of APs, and the payment that $v_{k}$ can receive, in order to jointly determine: (i) which ESs in $\mathcal{C}_{k}$ should be served without exceeding $v_{k}$’s remaining capacity $C_{k}^{\mathrm{A}}$, and (ii) in what visiting order these ESs are arranged along the route of $v_{k}$. By maximizing the AP’s utility over all such candidate routes, we obtain a selected ES set $\mathcal{K}_{k}^{\ast}\subseteq\mathcal{C}_{k}$ and the associated pre-planned path for AP $v_{k}$. Furthermore, for every $s_{i}\in\mathcal{K}_{k}^{\ast}$, we set $x_{i,k}^{\mathrm{Off}}=1$ and remove $s_{i}$ from $\mathcal{U}^{\mathrm{free}}$, ensuring that each ES is matched to (at most) one AP. Stacking these per-AP decisions over all $k$ yields the overall ES–AP assignment matrix $\bm{X}^{\ast}$. Moreover, since every successful pair $(s_{i},v_{k})$ satisfies $\overline{\mathrm{bid}}_{i}\geq\mathrm{ask}_{k}$, each winning ES is guaranteed not to pay more than its declared value in the matching stage, which contributes to ES-side individual rationality.
Step 3. Pricing and contract finalization via binary search: Once $\bm{X}^{\ast}$ is determined, Off-AIC2 computes ES payments and AP rewards through a monotone binary-search procedure that approximates critical-value pricing. The high-level goal is twofold: (i) discourage large deviations from truthful bidding, and (ii) keep the total ES payment and total AP reward closely aligned so that the mechanism does not incur a systematic deficit. In particular, for each matched ES buyer $s_{i}$ (with $\sum_{k}x^{\mathrm{Off}}_{i,k}=1$), we treat $\overline{\mathrm{bid}}_{i}$ as an effective bid parameter and search over the interval $\overline{\mathrm{bid}}_{I^{\ast}+1},\,\overline{\mathrm{bid}}_{i}]$ to find the smallest bid $p_{i}^{\mathrm{b}\ast}$ such that, if $s_{i}$ had declared this lower bid (while all other bids/asks remained unchanged) and Step 2 was re-executed, $s_{i}$ would still be selected as a winner. This $p_{i}^{\mathrm{b}\ast}$ is the ES’s critical payment: bidding any lower would cause $s_{i}$ to lose the contract. Symmetrically, for each matched AP seller $v_{k}$ (with $\sum_{i}x^{\mathrm{Off}}_{i,k}\geq 1$), we vary its effective ask within $\mathrm{ask}_{K^{\ast}+1},\,\mathrm{ask}_{k}]$ and re-run Step 2 to determine the largest reward level $r_{k}^{\mathrm{s}\ast}$ such that $v_{k}$ still remains in the winning set. This $r_{k}^{\mathrm{s}\ast}$ acts as a critical reward: if $v_{k}$ has a higher ask, it would not be matched to any ES.
The resulting settlement vectors $(\bm{p}^{\mathrm{b}\ast},\bm{r}^{\mathrm{s}\ast})$ finalize the offline contracts between ESs and APs under Off-AIC2. By tying payments and rewards to these critical values, the mechanism encourages near-truthful [^3] Unlike conventional auctions that ignore spatio-temporal aspects of the system, truthfulness in our setting is inherently shaped by spatial advantages: ESs close to AP routes can be profitably served with lower bids, whereas remote ESs incur higher travel and time costs and must bid higher to remain competitive; similarly, APs with route-aligned demand clusters face lower marginal costs than detour-intensive ones. As ES–AP matching is jointly coupled with route planning in eACO-VRP, strict dominant-strategy truthfulness cannot be guaranteed. Instead, the binary-search-based payment rule approximates critical-value pricing and induces near-truthful behavior in practice. reporting of private valuations while keeping buyer payments and seller rewards balanced at the system level (a property that we numerically demonstrate in Sec. VI-C3). In summary, Off-AIC2 satisfies individual rationality, approximate truthfulness, and budget balance. Since these properties follow standard definitions, we relegate their formal definitions and detailed proofs to Appx. B–E to maintain clarity and focus in the main text.
@@
V Online Stage: Potential Games for Service Scheduling under Human-Machine Coexistence
Unlike the offline stage, the online stage aims to cope with real-time service requests generated by HUs and MUs. In practice, the number and composition of SDs that actually arrive in each trading round may deviate from the offline forecasts. As a result, the offline outcome can only reserve a feasible pool of UAVs and service capacities for ESs with heavy resource demands, but it cannot directly determine which SP should serve which SD, at what time, or at what service rate given the real-time state of the system. Therefore, in this section we build on the offline contracts and capacity budgets, and design an online service scheduling mechanism that dynamically refines SD–SP associations and resource allocation so as to adapt to the current system state.
@@ V Online Stage: Potential Games for Service Scheduling under Human-Machine Coexistence
In particular, we model the online service scheduling through interactions among two players: (i) SPs (e.g., ESs and APs), collected as $\mathcal{S}^{\mathrm{On}}=\{s^{\mathrm{On}}_{1},\ldots,s^{\mathrm{On}}_{\mathbbm{i}},\ldots,s^{\mathrm{On}}_{|\mathcal{S}^{\mathrm{On}}|}\}$; and (ii) SDs (e.g., HUs and MUs), collected as $\mathcal{U}^{\mathrm{On}}=\{u^{\mathrm{On}}_{1},\ldots,u^{\mathrm{On}}_{\mathbbm{j}},\ldots,u^{\mathrm{On}}_{|\mathcal{U}^{\mathrm{On}}|}\}$. In this stage, each SD strategically decides whether to execute its tasks locally or offload them to one of the available SPs, while SPs allocate their computing capacities to their associated tasks. Such mutually coupled decisions are then considered as a potential game. Specifically, to represent the offloading decision of $u^{\mathrm{On}}_{\mathbbm{j}}$, we define the following binary indicator vector:
$\hskip-8.53581pt\bm{X}^{\mathrm{On}}_{\mathbbm{j}}=\big(x^{\mathrm{On}}_{\mathbbm{j},0},\,x^{\mathrm{On}}_{\mathbbm{j},1},\,\ldots,\,x^{\mathrm{On}}_{\mathbbm{j},\mathbbm{i}},\,\ldots,\,x^{\mathrm{On}}_{\mathbbm{j},|\mathcal{S}^{\mathrm{On}}|}\big)\in\{0,1\}^{\,1+|\mathcal{S}^{\mathrm{On}}|},$ (9)
where $x^{\mathrm{On}}_{\mathbbm{j},0}=1$ indicates local task execution at $u^{\mathrm{On}}_{\mathbbm{j}}$ and $x^{\mathrm{On}}_{\mathbbm{j},\mathbbm{i}}=1$ indicates offloading to SP $s^{\mathrm{On}}_{\mathbbm{i}}$ (note that $\sum_{\mathbbm{i}=0}^{|\mathcal{S}^{\mathrm{On}}|}x^{\mathrm{On}}_{\mathbbm{j},\mathbbm{i}}=1,~\forall\,\mathbbm{j}$). Given the profile $\bm{X}^{\mathrm{On}}_{\mathbbm{j}}$, the resulting delays and costs are jointly determined by how many SDs concurrently choose the same SP and how its computing resources are shared. We next formalize this joint effect through a delay and congestion model, which will serve as the basis for defining the utilities of SDs and SPs.
@@ V-A1 Delay and Congestion Model
We consider the SP set $\mathcal{S}^{\mathrm{On}}$ and the SD set $\mathcal{U}^{\mathrm{On}}=\bm{\mathcal{U}}^{\mathrm{H}}\cup\bm{\mathcal{U}}^{\mathrm{M}}$, where $\bm{\mathcal{U}}^{\mathrm{H}}$ and $\bm{\mathcal{U}}^{\mathrm{M}}$ denote the sets of HUs and MUs, respectively, where each SP $s^{\mathrm{On}}_{\mathbbm{i}}$ has a finite computing capacity $f^{\mathrm{On}}_{\mathbbm{i}}>0$ (CPU cycles/s). When a subset of SDs $\mathcal{U}_{\mathbbm{i}}=\{u^{\mathrm{On}}_{\mathbbm{v}}\in\mathcal{U}^{\mathrm{On}}:x^{\mathrm{On}}_{\mathbbm{v},\mathbbm{i}}=1\}$ simultaneously offload their workload to $s^{\mathrm{On}}_{\mathbbm{i}}$, the SP must multiplex its limited computing resource among heterogeneous requests with different priorities and contract terms. To capture both congestion effects and differentiated service levels in a tractable way, we adopt a weighted processor sharing (WPS) rule [13]: each SD $u^{\mathrm{On}}_{\mathbbm{j}}$ associated with $s^{\mathrm{On}}_{\mathbbm{i}}$ is assigned a positive weight $\gamma_{\mathbbm{j},\mathbbm{i}}>0$ that reflects its service priority, contract tier, or fairness quota. The effective computing rate that $u^{\mathrm{On}}_{\mathbbm{j}}$ receives from $s^{\mathrm{On}}_{\mathbbm{i}}$ is then given by $\hat{f}_{\mathbbm{j},\mathbbm{i}}=f^{\mathrm{On}}_{\mathbbm{i}}\frac{\gamma_{\mathbbm{j},\mathbbm{i}}}{\sum_{u^{\mathrm{On}}_{\mathbbm{v}}\in\mathcal{U}^{\mathrm{On}}_{\mathbbm{i}}}\gamma_{\mathbbm{v},\mathbbm{i}}}$, and the corresponding edge-side computing delay can be obtained as:
@@
where $r_{\mathbbm{j}}$ denotes the computing workload (e.g., CPU cycles needed) of $u^{\mathrm{On}}_{\mathbbm{j}}$. This WPS model compactly captures how the delay of each SD grows with the aggregate load at the SP. We further summarize the impact of associated SDs on $s^{\mathrm{On}}_{\mathbbm{i}}$ through an effective congestion load defined as $y_{\mathbbm{i}}=\sum_{u^{\mathrm{On}}_{\mathbbm{v}}\in\mathcal{U}^{\mathrm{On}}_{\mathbbm{i}}}\gamma_{\mathbbm{v},\mathbbm{i}}$. With this notation, (10) can be written in a factorized form as:
$t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}=\theta_{\mathbbm{j},\mathbbm{i}}\;\tau_{\mathbbm{i}}(y_{\mathbbm{i}}),\quad\theta_{\mathbbm{j},\mathbbm{i}}=\frac{r_{\mathbbm{j}}}{\gamma_{\mathbbm{j},\mathbbm{i}}},\quad\tau_{\mathbbm{i}}(y)=\frac{y}{f^{\mathrm{On}}_{\mathbbm{i}}},$ (11)
where $\theta_{\mathbbm{j},\mathbbm{i}}$ is a user-specific load factor, whereas $\tau_{\mathbbm{i}}(y)$ captures the server-side congestion effect as a function of the aggregate load. This separable structure will be instrumental for constructing an ordinal potential in the subsequent analysis.
While (11) characterizes the congestion-induced computing delay, an offloading SD also experiences wireless transmission latency in the uplink: when $u^{\mathrm{On}}_{\mathbbm{j}}$ offloads to $s^{\mathrm{On}}_{\mathbbm{i}}$, the uplink delay is $t^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}=\frac{d_{\mathbbm{j}}}{W\log_{2}\!\big(1+e^{\mathrm{tx}}_{\mathbbm{j}}\,\Gamma_{\mathbbm{j},\mathbbm{i}}\big)}$, where $d_{\mathbbm{j}}$ is the input data size, $W$ is the bandwidth, $e^{\mathrm{tx}}_{\mathbbm{j}}$ is the transmit power, and $\Gamma_{\mathbbm{j},\mathbbm{i}}$ is the Signal-to-Noise-Ratio (SNR) between $u^{\mathrm{On}}_{\mathbbm{j}}$ and $s^{\mathrm{On}}_{\mathbbm{i}}$. Consequently, the total edge-side latency experienced by $u^{\mathrm{On}}_{\mathbbm{j}}$ when uploading its workload to $s^{\mathrm{On}}_{\mathbbm{i}}$ is given by:
@@
V-A2 Utility of HUs
Building on the above latency and energy models, we now characterize the utility of an HU when deciding whether to execute a task locally or offload it to an SP. Consider an HU $u^{\mathrm{On}}_{\mathbbm{j}}\in\bm{\mathcal{U}}^{\mathrm{H}}$ with local CPU frequency $f^{\mathrm{loc}}_{\mathbbm{j}}$ and local computational power $e^{\mathrm{loc}}_{\mathbbm{j}}$. The local execution time and energy consumption of its task are given by $t^{\mathrm{loc}}_{\mathbbm{j}}={r_{\mathbbm{j}}}/{f^{\mathrm{loc}}_{\mathbbm{j}}}$ and $E^{\mathrm{loc}}_{\mathbbm{j}}={r_{\mathbbm{j}}\,e^{\mathrm{loc}}_{\mathbbm{j}}}/{f^{\mathrm{loc}}_{\mathbbm{j}}}$, respectively. When the HU $u^{\mathrm{On}}_{\mathbbm{j}}$ offloads its task to $s^{\mathrm{On}}_{\mathbbm{i}}$, it experiences latency $t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}+t^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}$ and uplink energy $E^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}$. Compared with local task execution, this leads to the latency saving of $t^{\mathrm{save}}_{\mathbbm{j},\mathbbm{i}}{=}t^{\mathrm{loc}}_{\mathbbm{j}}-\big(t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}+t^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}\big)$ and energy saving of $c^{\mathrm{save}}_{\mathbbm{j},\mathbbm{i}}{=}E^{\mathrm{loc}}_{\mathbbm{j}}-E^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}$ ($E^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}$ denotes the uplink transmission energy).
To map these latency and energy savings into a scalar benefit, we introduce non-negative valuation coefficients $\mathbb{V}^{\mathrm{H}}_{t}$ and $\mathbb{V}^{\mathrm{H}}_{e}$ to capture HU sensitivities to latency and energy, respectively. The HU’s perceived valuation of being served by $s^{\mathrm{On}}_{\mathbbm{i}}$ is then modeled via $v^{\mathrm{H}}_{\mathbbm{j},\mathbbm{i}}=\mathbb{V}^{\mathrm{H}}_{t}\,t^{\mathrm{save}}_{\mathbbm{j},\mathbbm{i}}+\mathbb{V}^{\mathrm{H}}_{e}\,c^{\mathrm{save}}_{\mathbbm{j},\mathbbm{i}}$. Given this valuation, the HU’s net utility must also account for the service payment: if $s^{\mathrm{On}}_{\mathbbm{i}}$ charges a price $p_{\mathbbm{j},\mathbbm{i}}$ for serving $u^{\mathrm{On}}_{\mathbbm{j}}$, its net utility is